Three students Camille, Danielle and Erika are in a race. Camille and Danielle have the same probability of winning is twice as likely to win as Erika. Find the probability that
a. Camille wins
b. Danielle wins
c. Erika wins

Answer by Edwin McCravy(20077)

x = P(E wins)
P(C wins) = 2x
P(D wins) = 2x

P(E wins OR C wins OR D wins) = x + 2x + 2x = 1
                                         5x = 1
                                          x = 1/5

P(E wins) = x  = 1/5
P(C wins) = 2x = 2(1/5) = 2/5
P(D wins) = 2x = 2(1/5) = 2/5 

Edwin